# Spectra for the Vacuum Cherenkov Effect in Astrophysical Electromagnetic Cascades with Lorentz Invariance Violation

### v1

@@ 1 Introduction
By allowing Lorentz symmetry to be broken, LIV leads to a wealth of new processes that would otherwise be forbidden, such as spontaneous photon decay, photon splitting, vacuum Cherenkov (VC) radiation from charged particles, and even spontaneous disintegration of atomic nuclei. This ultimately impacts how particles travel in the universe.
@@ 2 The kinematics of the vacuum Cherenkov effect
$E_{e}^{2}=m_{e}^{2}+p_{e}^{2}+\sum\limits_{n=0}^{\infty}\chi_{n}^{e}\dfrac{p_{e}^{n+2}}{M_{\rm Pl}^{n}}\;,\;\;\;E_{\gamma}^{2}=k_{\gamma}^{2}+\sum\limits_{n=0}^{\infty}\chi_{n}^{\gamma}\dfrac{k_{\gamma}^{n+2}}{M_{\rm Pl}^{n}}\,,$ (2)
@@ 3 Probabilities and Rates for the Vacuum Cherenkov Effect
$\begin{split}&\mathcal{G_{\pm}}\equiv\left[37\left(\pm\mathcal{S}-6\chi_{2}^{\gamma}\right)\chi_{2}^{e}\chi_{2}^{\gamma}-64\left(\chi_{2}^{e}\right)^{3}-\left(\pm 14\mathcal{S}-207\chi_{2}^{\gamma}\right)\left(\chi_{2}^{e}\right)^{2}-10\left(\pm 5\mathcal{S}-16\chi_{2}^{\gamma}\right)\left(\chi_{2}^{\gamma}\right)^{2}\right]\\ &\times\frac{\chi_{2}^{e}\left(\pm S-3\chi_{2}^{e}\right)}{160\left(\chi_{2}^{\gamma}-\chi_{2}^{e}\right)^{4}}\,,\;\;\;\text{and}\;\;\mathcal{G}_{0}\equiv\dfrac{157\chi_{2}^{e}-22\chi_{2}^{\gamma}}{120}\,,\end{split}$ (9)
wherein $\mathcal{S}\equiv\sqrt{3\chi_{2}^{\rm e}\left(4\chi_{2}^{\gamma}-\chi_{2}^{e}\right)}$. Employing these definitions, we can now understand the behavior of the VC rate considering the $\chi_{2}^{e}$$\chi_{2}^{\gamma}$ parameter space. The possible values of $\Gamma_{\rm VC}$ for all the different parameter combinations presented in Eq. (7).
@@
Based on the symmetry of the allowed values of $x$, cf. (7), we illustrate Eq. (10) by exemplarily considering parameter combinations which obey the relation $\left|\chi_{2}^{\gamma}\chi_{2}^{e}\right|=10^{-10}$.
@@ 4 Results
Proceeding to the electron spectra, for case A we see that the spectral shape of ${\rm d}N/{\rm d}E_{{\rm out},e}$ is a flat spectrum with a sharp cut-off which occurs when the electron momentum drops below the VC threshold value. The only spectral variation appears right below the cut-off where the spectrum either slightly rises or falls. This is, again, a direct consequence of the corresponding differential probability distribution function, but this time resulting in a reverse relationship compared to the photon spectra which is due to the fact that the electron carries away the remaining fraction of the incoming momentum, $1-x$.
@@ 5 Discussion and Outlook
While our primary focus was on electrons and positrons, our finding can be immediately extended to other charged leptons, namely muons and tauons, provided that their lifetime exceeds the VC time scale.

### v2

@@ 1 Introduction
By allowing Lorentz symmetry to be broken, LIV leads to a wealth of new processes that would otherwise be forbidden, such as spontaneous photon decay, photon splitting, vacuum Cherenkov (VC) radiation from charged particles, and even spontaneous disintegration of atomic nuclei [4]. This ultimately impacts how particles travel in the universe.
@@ 2 The kinematics of the vacuum Cherenkov effect
$E_{e}^{2}=m_{e}^{2}+p_{e}^{2}+\sum\limits_{n=0}^{\infty}\chi_{n}^{e}\dfrac{p_{e}^{n+2}}{M_{\rm Pl}^{n}}\;,\;\;\;E_{\gamma}^{2}=k_{\gamma}^{2}+\sum\limits_{n=0}^{\infty}\chi_{n}^{\gamma}\dfrac{k_{\gamma}^{n+2}}{M_{\rm Pl}^{n}}$ (2)
@@ 3 Probabilities and Rates for the Vacuum Cherenkov Effect
$\Gamma_{\rm VC}=\begin{cases}\Gamma_{1}\equiv\alpha\mathcal{G}_{0}\dfrac{p_{{\rm in},e}^{3}}{M_{\rm Pl}^{2}}\,,&x\in(0;1)\,,(\chi_{2}^{e}\geq 0)\wedge(\chi_{2}^{\gamma}\leq\chi_{2}^{e})\,,\\ \Gamma_{2}\equiv\alpha\mathcal{G}_{+}\dfrac{p_{{\rm in},e}^{3}}{M_{\rm Pl}^{2}}\,,&x\in(0;x_{\rm VC,+})\,,0<\chi_{2}^{e}<\chi_{2}^{\gamma}\,,\\ \Gamma_{3}\equiv\alpha(\mathcal{G}_{0}-\mathcal{G}_{-})\dfrac{p_{{\rm in},e}^{3}}{M_{\rm Pl}^{2}}\,,&x\in(x_{\rm VC,+};1)\,,\chi_{2}^{\gamma}<\chi_{2}^{e}<0\,,\\ [\text{no VC}]\,,&\text{else,}\end{cases}$ (5)
@@
[…]
wherein $\mathcal{S}\equiv\sqrt{3\chi_{2}^{\rm e}\left(4\chi_{2}^{\gamma}-\chi_{2}^{e}\right)}$. Employing these definitions, we can now understand the behavior of the VC rate considering the $\chi_{2}^{e}$$\chi_{2}^{\gamma}$ parameter space. The possible values of $\Gamma_{\rm VC}$ for all the different parameter combinations are presented in Eq. (7).
@@
Based on the symmetry of the allowed values of $x$, cf. Eq. (7), we illustrate Eq. (10) by exemplarily considering parameter combinations which obey the relation $\left|\chi_{2}^{\gamma}\chi_{2}^{e}\right|=10^{-10}$.
@@ 4 Results
Proceeding to the electron spectra, for case A we see that the spectral shape of ${\rm d}N/{\rm d}E_{{\rm out},e}$ is a flat spectrum with a sharp cut-off which occurs when the electron momentum drops below the VC threshold value. The only spectral variation appears right below the cut-off where the spectrum either slightly rises or falls. This is, again, a direct consequence of the corresponding differential probability distribution function, but this time resulting in a reverse relationship compared to the photon spectra which is due to the fact that the electron carries away the remaining fraction of the incoming momentum, 1x1-x.
@@ 5 Discussion and Outlook
While our primary focus was on electrons and positrons, our findings can be immediately extended to other charged leptons, namely muons and tauons, provided that their lifetime exceeds the VC time scale.